a 43 foot long string is cut into 4 pices the first piece of string is twice as long as the second piece of string. the 3rd piece of string is 3 feet longer than the second pice of string the fouth piece is twice as long as the first piece.what is the length of thr longest piece of string/
Answer by edjones(8007)

Let the 1st piece=a, 2nd=b, 3rd=c and the 4th=d
a=2b
c=b+3
d=2a=2(2b)=4b
2b+b+b+3+4b=43
8b+3=43
8b=40
b=5
a=10
c=8
d=20
.
Ed